1. Determine the peak frequency deviation for a FM modulator with a deviation sensitivity of 12 kHz/V  and a peak modulating signal of 9V.
1
Expert's answer
2021-10-26T17:15:49-0400

We know that

Voltage "V=V_0sin(wt)"

"V=9sin(24\pi t+\Phi)"

Peak frequency deviation

"n=\frac{12}{\sqrt2}=6\sqrt{2}Hz"